OpenStudy (anonymous):

x^2-3x-54=0 Can someone show me how to work this out?

OpenStudy (anonymous):

factor (x-9) (x+6) = 0 equate x-9 = 0 and x+6 = 0 x= 9 and x = -6

OpenStudy (anonymous):

That makes better sence to me thanks